OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4).

OpenStudy (anonymous):

\[Poin-Slope \space Form = y - y_1 = m(x - x_1)\]where m = slope and \((x_1, y_1)\)

OpenStudy (anonymous):

I dont understand slope intercept at all

OpenStudy (anonymous):

THis isn't slope-intercept though...

OpenStudy (anonymous):

Its slope something

OpenStudy (anonymous):

Just plug the numbers in... m = -3 \(x_1\) = 3 \(y_1\) = -4

OpenStudy (anonymous):

Is that seriously all I had to do? Why couldnt the teacher of just said that?!?!

OpenStudy (anonymous):

You just have to know where the numbers come from.

OpenStudy (anonymous):

So if thats how you do that. How do you do this? Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4).

OpenStudy (anonymous):

You have to first find the slope. Do you know the formula to find it?

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

THat's the form for slope-intercept. That comes later. The slope formula is: \[m = \frac{y_1 - y_2}{x_1 - x_2}\]

OpenStudy (anonymous):

So, you have two coordinates. (2, 4) and (5, 4) Now, let's say that (2, 4) is \((x_1, y_1)\). Does it make sense so far? That would make \(x_1 = 2\) and \(y_1 = 4\)

OpenStudy (anonymous):

Could you tell me what \(x_2\) and \(y_2\) are?

OpenStudy (anonymous):

2 and 4?